### 11 Dec 2023 v1

# QuickQuakeBuildings: Post-earthquake SAR-Optical Dataset for Quick Damaged-building Detection

Abstract
This is the preprint version. Quick and automated earthquake-damaged building detection from post-event satellite imagery is crucial, yet it is challenging due to the scarcity of training data required for developing robust algorithms. This letter presents the first dataset dedicated to detecting earthquake-damaged buildings from post-event very high resolution (VHR) Synthetic Aperture Radar (SAR) and optical imagery. Utilizing open satellite imagery and annotations acquired after the 2023 Turkey–Syria earthquakes, we deliver a dataset of co-registered building footprints and satellite image patches of both SAR and optical data, encompassing more than four thousand buildings. The task of damaged building detection is formulated as a binary image classification problem, that can also be treated as an anomaly detection problem due to extreme class imbalance. We provide baseline methods and results to serve as references for comparison. Researchers can utilize this dataset to expedite algorithm development, facilitating the rapid detection of damaged buildings in response to future events. The dataset and codes together with detailed explanations are made publicly available at https://github.com/ya0-sun/PostEQ-SARopt-BuildingDamage.
Index Terms:
building damage detection, convolutional neural network (CNN), very high resolution (VHR), remote sensing imagery, synthetic aperture radar (SAR), earthquake, geographic information system (GIS), OpenStreetMap (OSM), large-scale urban areas.
I Introduction
Earthquakes can result in substantial structural and infrastructural damage, often with significant socioeconomic consequences. After an event, fast and accurate detection of earthquake-damaged buildings in remote sensing imagery is of great importance. Remote sensing technologies can effectively improve the efficiency of disaster management and have been employed to estimate the extent of earthquake damage to buildings [1, 2, 3]. Very high resolution (VHR) optical images are easier to interpret, making them a preferred choice for many studies, if available. SAR imagery, on the other hand, is particularly interesting due to its capability of acquiring information regardless of weather and sun illumination conditions. Moreover, the enhanced resolution of contemporary SAR satellite images enables the extraction of information at the individual building level [4, 5, 6], comparable to VHR optical data.
Over the past years, many researchers have developed algorithms utilizing SAR data to detect earthquake-damaged buildings. Most existing works detect changes using both pre- and post-earthquake SAR images to acquire building damage information [7, 8]; however, pre-event high-resolution SAR imagery is generally unavailable in most locations. A few works simulate a SAR image using building shapes extracted from a pre-event optical image and acquisition parameters of a post-event SAR image and detect earthquake-damaged buildings by comparing the simulated and real SAR images  [9]. Utilizing only a single post-event SAR image, some researchers detect damaged buildings by analyzing signatures of destroyed buildings in high-resolution SAR data [10, 11, 12]. However, the study areas often comprise a limited number of isolated buildings, which do not accurately depict the typical conditions in densely populated urban areas. Therefore, the question of whether a single VHR SAR image acquired after an event can effectively identify damaged buildings remains to be addressed. In this regard, benchmark datasets play a pivotal role in the development and comparative assessment of diverse methodologies aimed at addressing the following questions: To what extent can a single VHR SAR image, acquired post-event, allow to identify damaged buildings and with which accuracy? Additionally, how do the outcomes derived from a single post-event SAR image compare with those of an optical image?
From a practical standpoint, there are several challenges in creating such a dataset: 1) Limited availability of high-resolution SAR images in disaster-affected areas; 2) Absence of labels for damaged buildings; 3) Lack of accurate terrain models, without which aligning the two is a complex task. Currently, such a dataset does not exist in the remote sensing field.
This work presents the first dataset for detecting earthquake-damaged buildings in post-event SAR and optical satellite imagery. We integrate publicly accessible satellite imagery and annotations obtained following the 2023 Turkey-Syria earthquakes and construct a dataset comprising over four thousand buildings, each with satellite image patches of both post-event SAR and optical data and its footprint co-registered with the corresponding image patches. We formulate the problem of damaged building detection as an image classification task and benchmark a set of popular machine learning and deep learning methods as a baseline reference.
The remaining part of the letter proceeds as follows: Section II introduces the dataset generation approaches, Section III presents the baseline methods and results, and Section IV concludes the paper.
II Dataset Generation
This work aims at addressing building-level damages. The tasks of identifying pixel-level damages and classifying the types of damages are beyond the scope of this dataset.
Considering the increasing availability of building footprint information across various geographic locations, we integrate pre-event building footprints with post-event satellite images and assess whether buildings at those locations are damaged. The problem of detecting damaged buildings in post-event imagery is therefore formulated as an image classification task with two classes: damaged and intact buildings. We generate and deliver a dataset comprising post-event VHR SAR and optical image patches for each building, building footprint masks corresponding to the image patches, and labels indicating whether each building is damaged or intact. The image patches and masks serve as inputs for the algorithm, while the labels represent the ground truth.
II-A Study Area and Data Sources
The study area is chosen in the city of Islahiye, located in southeastern Turkey near the northwestern border of Syria. On February 6, 2023, a devastating magnitude 7.8 earthquake struck the Turkish province of Kahramanmaras, followed by a 7.5 magnitude aftershock nine hours later. The earthquake inflicted widespread destruction, resulting in a significant loss of life, numerous injuries, and extensive damage to buildings and infrastructure. Islahiye was one of the most affected areas.
After the earthquakes, a set of high-resolution satellite data was released under CC BY 4.0 license [^1] https://creativecommons.org/licenses/by/4.0/ to support rescue operations by commercial remote sensing companies, such as Maxar and Planet Lab for optical data, and Capella Space in the SAR domain. During the humanitarian relief efforts, communities across the globe, such as OSM and UN mappers, organized labeling events and identified and validated a significant number of damaged buildings.
We utilize the Spotlight SAR image from Capella Space, acquired on February 9, 2023. The SAR image is of type Geocoded Terrain Corrected (GEO) [^2] https://support.capellaspace.com/hc/en-us/articles/360039702691-SAR-Data-Formats, with a pixel spacing of 0.35 m in both the azimuth and the range direction. The incidence angle of this SAR image is ${}^{\circ}$. Figure 1 shows the image coverage and zoomed-in views of selected areas. The optical image covering the same area was obtained from Maxar Analysis-Ready Data (ARD) under Maxar’s open data program. The image was acquired on February 7, 2023 by WorldView-3, with a ground sampling distance of 0.31 m and an incidence angle of ${}^{\circ}$. In addition, we obtained post-event building footprints and labels of destroyed buildings in the study area from Humanitarian OpenStreetMap Team  [^3] https://data.humdata.org/dataset/hotosm_tur_buildings  [^4] https://data.humdata.org/dataset/hotosm_tur_destroyed_buildings. Since we consider both SAR and optical images, all used data are chosen or projected to the Universal Transverse Mercator (UTM) coordinate system so that they can be processed uniformly. The SAR image is logarithmically scaled in dB for further processing.
Figure 1: The SAR image coverage and zoomed-in views of three areas in the colored boxes in the SAR image, respectively.
                 before registration
                 after registration
Figure 2: Examples of four study areas in the SAR image. Building footprint polygons before and after registration are plotted in red and green, respectively.
II-B Co-registration of building footprints and satellite imagery
Building-level analysis requires accurate registration of 2-D building footprints with satellite images. The ARD optical image aligns well with building footprints, requiring no additional registration. For the GEO SAR image, inspection shows that building polygons are not well-matched with the SAR image, as shown in the first row of Figure 2, and further registration is needed.
In urban areas, the geocoding errors in SAR data are often caused by inaccurate terrain heights, as illustrated in Figure 3(a). A height error $\delta H$ causes an error of $\delta L$ in the slant range and a shift of $\delta G$ on the ground. For the used SAR image, the incidence angle is ${}^{\circ}$; thus, a height error of 10 meters results in an error of 10.69 meters on the ground, causing errors of $\delta x$ and $\delta y$ in the geocoded image related to the flight direction, as shown in Figure 3(b). The height error $\delta H$ is usually inconstant over the observed area by the SAR sensor; hence, so are the geocoding errors.
To improve the alignment of building polygons and the SAR image, we apply the algorithm developed in [13], which relies on the corresponding building features representing the bottom of sensor-visible walls in both the two data, i.e., double bounce lines in the SAR image and near-range boundaries of 2-D building polygons, as illustrated in Figure 4. As the majority of buildings remain upright, with expected double bounce line signatures on the SAR image, the algorithm is applicable. Interested readers are referred to  [13] for more details. Next, we briefly explain the main steps of the algorithms.
Figure 3: The geocoding error from inaccurate height. (a) $H_{t}$ and $H_{f}$ are the accurate height and inaccurate height of a point, and $\theta$ is the incidence angle. The height error $\delta H$ results in an error of $\delta L=\delta Hcos\theta$ in the slant range and $\delta G=\delta Hcot\theta$ on the ground. (b) In the geocoded image, $\delta G$ is decomposed to $\delta x$ and $\delta y$ in the image coordinate system.
Figure 4: Building signature in SAR imagery and building’s geometric correspondence between SAR and building polygon: the near-range side of the building footprint corresponds to the double bounce line in the SAR image, which is the far-range side of the facade signatures [13].
II-B1 Extracting corresponding features
In the SAR image, double bounce lines correspond to the far-range side of the bright building signature. The SAR image is first segmented using Potts model [14]. Then, an intensity threshold and an area threshold are applied to select building wall segments. Subsequently, the boundary of wall segments is extracted, and the visibility check is performed to extract the double bounce lines caused by the wall segment boundary.
In the building polygons, the segments in the near-rage side of building polygons are extracted from each footprint, which in 3-D represent the bottom of illuminated or partially illuminated walls.
II-B2 Registering corresponding features
two point sets, GIS points (from building polygons) and SAR points, are then sampled in the extracted features from both data, and the registration problem is reduced to determine the correspondence and the underlying spatial transformation between two point sets.
Feature registration consists of three progressive steps: global registration, subarea registration, and polygon registration, and the rigid registration in each step is solved with the Iterative Closest Point (ICP) algorithm [15]. Global registration uses rough height values for an initial alignment of the two data, ensuring that the residual shift falls within a manageable range. Then, a set of grids, i.e., subareas, is evenly distributed over the whole region. The distance between one GIS point and its closest SAR point is calculated for all points. If the distribution of all distances within one subarea shows a clear center, the $\delta H$ in each grid is considered to be constant, and subarea registration is conducted. When the distribution of all distances does not show a clear center, the constant $\delta H$ assumption does not hold, and the registration proceeds to the polygon level, i.e., finding a rigid transformation for each polygon.
Since the SAR image we used, i.e., the GEO product, has been terrain corrected, we neglect global registration and perform subarea and polygon registration. The direction of transformation, GIS to SAR or SAR to GIS, is inconsequential for later analysis, and thus, we add the transformations of each step to building polygons. Afterward, manual validation and editing with expert knowledge are conducted to ensure accurate registration within the study area. Figure 2 shows building polygons on the SAR image in exemplary areas before and after registration. As can be seen, the registration procedure effectively aligns the two data sets. Note that buildings located around the mountainous area on the west side of the city, e.g., the area in and near the red box shown in Figure 1, are discarded due to unclear signals in the SAR image, which makes it difficult to check registration results.
II-C Patch generation
a b c d e f
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Examples of the dataset: a, b, c are intact buildings, and d, e, f are damaged buildings.
For each building, we crop the SAR image and the optical image based on the area of the building, considering including the target building area, i.e., footprint, wall, and roof, as well as the possible ruins around the target building and excluding surrounding buildings.
For side-looking SAR images, layover areas of buildings extend from building footprints towards the near-range direction. Therefore, when cropping SAR image patches, a buffer size of 10 pixels, i.e., around 3.5 m, is counted for far-range sides, and for near-range sides, an additional buffer size is counted to include layover areas of the target building in the image patch. Specifically, the layover length corresponding to 50 m building height in the ground range direction is decomposed to the image x- and y- directions and added to the buffer size. The optical image has a small off-nadir angle of ${}^{\circ}$, resulting in a minor offset between the roof outline and the corresponding 2-D building footprint when the image is not perfectly orthorectified. To ensure that the image patches include the entire building roof, we apply a 16-pixel buffer to compensate for the offset, which is approximately 5 meters. This buffer is slightly larger than the one used when cropping the SAR image from the near-range side, and it is used to crop the optical image around the bounding box of the building footprint polygons.
In addition, for each building, its footprint mask is generated corresponding to the SAR patch and the optical patch, respectively. For side-looking SAR data, it is necessary to include the footprint mask to help locate the target building, as the SAR patch may include signals of surrounding buildings. For nadir-looking optical data, footprint masks are included for a fair comparison with the SAR data.
Among the 4029 buildings in the study area, 169 are damaged, and the remaining 3860 buildings are intact. Consequently, the dataset contains 169 damaged buildings and 3860 intact buildings, and each of them has four patches: a SAR image patch, a SAR footprint patch, an optical image patch, and an optical footprint patch. Figure 5 shows examples of the dataset on six buildings, of which three are intact and three are damaged.
III Experimental Results
We benchmark an image classification task of two classes: damaged buildings and intact buildings. As the two classes are highly imbalanced between damaged and intact buildings, it can also be considered as an anomaly detection task, i.e., detecting damaged buildings in the whole dataset.
III-A Baseline Approaches
Four methods are introduced to evaluate the performance and to establish a benchmark: support vector machine (SVM), random forests (RF), a 3-layer convolutional neural networks (CNN), and ResNet-18 [16].
SVM and RF are selected for their good performance in many applications, including classifying collapsed and standing buildings from post-event SAR imagery as reported in [11]. To extract features as the input, we follow the experimental setup proposed in [11], which employs four first-order statistics, i.e., mean, variance, skewness, and kurtosis, and eight second-order image statistical measures, i.e., mean, variance, homogeneity, contrast, dissimilarity, entropy, second moment, and correlation. For a detailed explanation, the readers are referred to [11].
A simple 3-layer CNN and a ResNet-18 are selected as the two deep learning methods. The simple CNN consists of 3 convolution-ReLU-maxpool blocks, followed by average pooling, a linear layer with dropout, and a final classification layer. The ResNet-18 follows the standard design in [16], consisting of 18 convolutional layers with residual connections. Similar to the simple CNN, we also add a linear layer with dropout before the final classification layer.
III-B Evaluation Metrics
To evaluate the performance of the baseline methods, we report the precision, recall, and $F_{1}$ scores:
$P=\frac{tp}{tp+fp},R=\frac{tp}{tp+fn},F_{1}=2\cdot\frac{P\cdot R}{P+R},$ (1)
where $P$ and $R$ denote the precision and recall, respectively, and $tp$, $fp$, $tn$, $fn$ represent true positives, false positives, true negatives, and false negatives for buildings, respectively.
In addition, we report the area under the receiver-operator curve (AUROC), a standard metric in anomaly detection tasks that well reflects the model’s efficiency in distinguishing between classes. The AUROC score summarizes the ROC curve into a single number that describes the performance of a model for multiple thresholds at the same time.
III-C Implementation Details
We conduct cross-fold experiments for a robust evaluation on the relatively small dataset. Specifically, we split the dataset into 5 folds with a balanced number of damaged and intact buildings, and run each experiment 5 times with 4 folds training and 1 fold testing. The mean and standard deviation are calculated and reported for every evaluation metric.
We preprocess the images by removing 2% pixel outliers and normalizing the pixel values to the range [0,1]. For simple CNN, we randomly initialize the model; for ResNet-18, we use ImageNet pretrained weights. We use random resized crop and random horizontal and vertical flip as data augmentations. To deal with the significant class imbalance, we give a bigger weight to damaged buildings and a smaller weight to intact buildings during data sampling. We optimize binary cross entropy loss with AdamW optimizer for 30 epochs. The learning rate follows a cosine-decay schedule starting from 0.0001. Batch size is set to 32.
For the statistic features used by the two machine learning methods, we calculated the features of image patches based on the implementation of or PyFeats [^5] https://github.com/giakou4/pyfeats. For calculating the evaluation metrics, we use the implementation in scikit-learn [^6] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.roc_auc_score.html.
III-D Performance Comparison
Table I illustrates variations in performance across different models on the dataset. For the SAR image, the SVM model shows the lowest precision but excels in recall (0.495), indicating it detects a high proportion of true positives but also has many false positives. RF outperforms other models in precision but has a lower recall. While having moderate precision and recall, CNN produces a relatively high $F_{1}$ score and the highest AUROC of 0.739 among SAR models, indicating its proficiency in handling complex features directly from raw SAR data. ResNet-18 underperforms CNN regarding all four metrics despite its higher complexity.
For the optical image, ResNet-18 stands out with the highest precision, recall, $F_{1}$ score, and AUROC. It outperforms other models across all metrics. The other deep learning model, CNN, demonstrates a balance between precision and recall (0.432), resulting in a high $F_{1}$ score and an impressive AUROC of 0.853. RF attains high precision at 0.611 but a relatively lower recall rate. In contrast, SVM yields less favorable outcomes compared to other models. Nevertheless, it is noteworthy that SVM’s AUROC of 0.723 surpasses that of three other models in the SAR image category.
Comparing SAR and optical images, we can see that SAR images appear to be more challenging for all models, with generally lower performance than optical images. For the two deep learning models, we observe that ResNet-18 excels in the optical images; however, CNN outperforms ResNet-18 regarding SAR images. The simple CNN’s better performance on SAR data might be attributed to several factors. SAR data’s unique characteristics may favor simpler models that efficiently capture essential patterns. The risk of overfitting is reduced with simpler models, which can lead to better generalization, particularly in cases of limited training data. In contrast, the complex deep neural networks may not be optimized for SAR-specific traits.
TABLE I: Benchmark results on the dataset: 5-fold mean (std). The highest values of different metrics are highlighted in bold.
image model Precision Recall $F_{1}$ AUROC
SAR SVM 0.096 (0.021) 0.495 (0.202) 0.154 (0.037) 0.653 (0.068)
RF 0.223 (0.090) 0.344 (0.208) 0.231 (0.071) 0.670 (0.079)
CNN 0.167 (0.028) 0.371 (0.073) 0.228 (0.032) 0.739 (0.026)
ResNet-18 0.155 (0.067) 0.276 (0.080) 0.184(0.050) 0.653 (0.043)
Optical SVM 0.163 (0.044) 0.352 (0.148) 0.207 (0.057) 0.723 (0.047)
RF 0.611 (0.215) 0.354 (0.098) 0.421 (0.106) 0.810 (0.062)
CNN 0.489 (0.154) 0.432 (0.078) 0.449 (0.096) 0.853 (0.045)
ResNet-18 0.666 (0.121) 0.539 (0.147) 0.581 (0.100) 0.938 (0.022)
IV Conclusion
Detecting earthquake-damaged buildings in post-event satellite imagery is essential yet challenging. This study introduced a dataset designed to address the issue of data scarcity, a crucial factor in fostering the development of robust algorithms for this purpose. The dataset combines post-event SAR and optical satellite images with labels of damaged and intact buildings, and the problem is formulated as an image classification task. We provided a benchmark on both modalities with four baseline methods. Results show that detecting damages from post-event SAR images is valuable and possible but more challenging than optical images. Such findings call for further research on improved methods, in particular on SAR images.
Constrained by limited data quantity and data imbalance, this dataset serves as a starting point. The dataset will undergo expansion and updates as new data emerges in the future. We hope that the research community will engage in further algorithm development for post-earthquake damaged building assessment in SAR images and, where feasible, share their data, thus expediting the identification of post-disaster damaged structures.
We refrained from data fusion as optical and SAR data are typically not simultaneously acquired. The dataset primarily aims to assess individual post-event SAR image performance and compare it with optical images. Nevertheless, the fusion of multi-modal data for post-earthquake building damage assessment may potentially enhance the overall effectiveness, offering a more comprehensive analysis of the damaged buildings when the data are available.
Acknowledgment
The authors would like to thank Capella Space (https://www.capellaspace.com/) and Maxar Technologies (https://www.maxar.com/) for providing satellite imagery and OpenStreetMap and Humanitarian OpenStreetMap Team (https://www.hotosm.org/) for providing building footprints and annotations of destroyed buildings.